# Wukong: Towards a Scaling Law for Large-Scale Recommendation

### v3

@@ Abstract
Scaling laws play an instrumental role in the sustainable improvement in model quality. Unfortunately, recommendation models to date do not exhibit such laws similar to those observed in the domain of large language models, due to the inefficiencies of their upscaling mechanisms. This limitation poses significant challenges in adapting these models to increasingly more complex real-world datasets.
In this paper, we propose an effective network architecture based purely on stacked factorization machines, and a synergistic upscaling strategy, collectively dubbed Wukong, to establish a scaling law in the domain of recommendation. Wukong’s unique design makes it possible to capture diverse, any-order of interactions simply through taller and wider layers.
We conducted extensive evaluations on six public datasets, and our results demonstrate that Wukong consistently outperforms state-of-the-art models quality-wise. Further, we assessed Wukong’s scalability on an internal, large-scale dataset. The results show that Wukong retains its superiority in quality over state-of-the-art models, while holding the scaling law across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example or equivalently up to Large Launguage Model (GPT-3) training compute scale, where prior arts fall short.
@@
Figure 1: Wukong outperforms existing state-of-the-art models while demonstrating a scaling law in the recommendation domain across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example or equivalently up to Large Launguage Model (GPT-3) training compute scale, where prior arts fall short.
@@ 1 Introduction
From movie recommendations to feed compilations, deep learning-based recommendation systems (DLRS) are powering a wide range of online services as reported by major content providers (Naumov et al., 2019; Wang et al., 2021a; Lian et al., 2021; Liu et al., 2022; Covington et al., 2016).
@@
To date, the primary trend of DLRS up-scaling is through sparse scaling, i.e., expanding the sizes of embedding tables by adding more entries to reduce collision and widening their dimensions for increased expressiveness. Consequently, DLRS have reached trillions of parameters (Kang et al., 2020; Mudigere et al., 2021; Lian et al., 2021; Luo et al., 2024) with embedding tables dominating the parameter count.
@@
Unfortunately, the traditional way of up-scaling has a few practical drawbacks. Merely expanding the sparse component of a model does not enhance its ability to capture the complex interactions among an increasing number of features. Moreover, this trend notably diverges from the trend of hardware advancements, as most improvements in the next generation accelerators lie in the compute capacity (Luo et al., 2018, 2017), which embedding table lookups cannot utilize. Thus, simply expanding embedding table leads to prohibitive infrastructure costs with suboptimal accelerator utilization, especially in distributed settings (Luo et al., 2024).
@@
Our work aims to find an alternative scaling mechanism for recommendation models, that can establish the scaling law, similar to that established in the LLM domain. Namely, we would like to devise a unified architecture whose quality can be continuously improved in conjunction with dataset size, compute and parameter budgets, with a synergistic strategy.
@@
We focus on upscaling interaction components, dubbed dense scaling, to mitigate the quality and efficiency drawbacks from sparse scaling. However, existing models cannot benefit from this paradigm for various reasons. For example, DLRM faces scalability constraints, being limited to capturing only second-order interactions, which restricts its effectiveness on complex datasets. Moreover, existing models lack a comprehensive process for scaling up. For example, DCNv2 and AutoInt+ focus on tuning certain hyperparmeters, leading to rapidly diminishing returns when scaling up. Additionally, even with modern tricks like residual connection (He et al., 2016), layernorm (Ba et al., 2016), gradient clip (Pascanu et al., 2013)), up-scaling existing models is prone to training stability issues (Tang et al., 2023).
@@
To establish a scaling law for recommendation models, we propose Wukong, a simple yet effective interaction architecture that exhibits effective dense scaling properties. Inspired by the principles of binary exponentiation, our key innovation is to use a series of stacked Factorization Machines (FMs) to efficiently and scalably capture any-order feature interactions. In our design, each FM is responsible of capturing second order interactions with respect to its inputs, and the outputs from these FMs are subsequently transformed by MLPs into new embeddings, which encode the interactions results and serve as inputs to the next layers.
@@
We evaluated Wukong’s performance using six public datasets and a large-scale internal dataset. The results demonstrate that Wukong outperforms state-of-the-art models across all public datasets in terms of AUC, indicating the effectiveness of Wukong’s architecture and its ability to generalize across a wide range of recommendation tasks and datasets. On our internal dataset, Wukong not only significantly outperforms existing models in terms of quality at comparable levels of complexity but also shows continuous enhancements in quality when scaled up across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example or equivalently up to 2336 PF-days of total training compute, normalized to 365 days assuming online training (Zhai et al., 2024), comparable to Large Launguage Model (GPT-3, 3640PF-days) training compute scale (Brown et al., 2020), where prior arts fall short.
@@ 2 Related Work
Dense Interaction Architectures    Capturing interaction between features is the key to effective DLRS, and various prior arts have been proposed for this cause. For instance, AFN+ (Cheng et al., 2020) transforms features into a logarithmic space to capture arbitrary order of interactions; AutoInt+ (Song et al., 2019) uses multi-head self-attention; DLRM and DeepFM (Naumov et al., 2019; Guo et al., 2017) leverage Factorization Machines (FM) (Rendle, 2010) to explicitly capture second order interactions; HOFM (Blondel et al., 2016) optimizes FM to efficiently capture higher order of interactions; DCNv2 (Wang et al., 2021a) uses CrossNet, which captures interactions via stacked feature crossing, which can be viewed as a form of elementwise input attention. FinalMLP (Mao et al., 2023) employs a bilinear fusion to aggregate results from two MLP streams, each takes stream-specific gated features as input. MaskNet (Wang et al., 2021b) adopts a series of MaskBlocks for interaction capture, applying “input attention” to the input itself and intermediate activations of DNN; xDeepFM (Lian et al., 2018) combines a DNN with a Compressed Interaction Network, which captures interactions through outer products and compressing the results with element-wise summation.
Scaling up DLRS    Existing literature mainly focuses on scaling up the sparse component of models (Kang et al., 2020; Mudigere et al., 2021; Lian et al., 2021). There is limited discussion on the scalability of dense interaction architectures. Among the few investigated, their scalability are often unsatisfactory. For example, AutoInt+ observes diminishing gains with more than two layers on public datasets. Similarly, DCNv2 experiences rapidly diminishing returns when expanding beyond two layers or increasing the rank size, despite being tested on a large-scale production dataset.
@@ 3.2 Embedding Layer
Given a multi-hot categorical input, an embedding table maps it to a dense embedding. This process involves a series of lookups, each corresponding to a “hot” dimensions within the input. The lookup results are then aggregated using a pooling operation, with summation being the typical method in the realm of embeddings.
In our design, the embedding dimension is standardized for all embeddings generated by the Embedding Layer, known as the global embedding dimension $d$. To accommodate the varying significance of different features, multiple embeddings are generated for each feature deemed significant. In contrast, less important features are allocated smaller underlying embedding dimensions. These smaller embeddings are then collectively grouped, concatenated, and transformed into $d$-dimensional embeddings using a MLP.
@@ 3.3 Interaction Stack
The interaction modules stack $l$ identical interaction layers, where each layer captures progressively higher-order feature interactions at an exponential rate using Factorization Machines (FMs).
@@ 3.4 Factorization Machine Block (FMB)
A FMB contains a Factorization Machine (FM) followed by a MLP. The FM is used to capture explicit feature interactions of the input embeddings, with the output being a 2D interaction matrix where each element represents the interaction between a pair of embeddings. This interaction matrix is flattened and converted to a vector with shape of $(n_{F}\times d)$ through the MLP, and reshaped to $n_{F}$ embeddings for later use.
@@ 3.6 Optimized FM
In its basic form where a pairwise dot-product is used, FM’s computation and storage complexity grows quadratically with the number of embeddings. This quickly becomes prohibitive on real-world datasets with thousands of features.
@@ 3.8 Scaling Wukong
Having defined the Wukong architecture, we summarize the main hyperparameters that are related to scale up and later we describe our efforts to upscaling Wukong with respect to these hyperparameters.
@@ 3.9 Intuition Behind Wukong’s Enhanced Effectiveness
Additionally, Wukong treats each embedding as a single unit, focusing on embedding-wise interactions. This approach significantly reduces computational demands compared to architectures that capture element-wise interactions, as it avoids the less useful intra-embedding and cross-dimensional interactions. Consequently, Wukong not only enhances the efficiency of the recommendation system but also maintains its effectiveness in capturing relevant feature interactions.
@@ 4 Implementation
Overall, distributed training is required to make Wukong training feasible. For the embedding layer, we use a column-wise sharded embedding bag implementation provided by Neo (Mudigere et al., 2021) and NeuroShard (Zha et al., 2023). On the dense part, we balance the trade-off between performance and memory capacity by adopting FSDP (Zhao et al., 2023) and tune the sharding factor so that the model fits in the memory without creating too much redundancy.
To enhance training efficiency, we employ both automatic operator fusion through torch.fx (Reed et al., 2022) to improve training performance. In addition, we aggressively apply quantization to reduce compute, memory, and communication overheads simultaneously. Specifically, we train Wukong’s embedding tables in FP16, and communicate embedding lookup results in FP16 in the forward pass and BF16 in the backward pass; we use BF16 quantization during the transport and reduction of gradients for dense parameters in the backward pass.
@@ 6.1.3 Metrics
AUC    Area Under the Curve (AUC) provides an aggregated measure of the model’s ability to correctly classify positives and negatives across all thresholds. A higher AUC value indicates a better model performance.
LogLoss    The log loss quantifies the penalty based on how far the prediction is from the actual label. The lower the log loss, the better the model.
@@ 7 Evaluation on an Internal Dataset
In this section, our demonstrate the scalability of Wukong and gain a deep understanding of how different individual components of Wukong contribute to its effectiveness, using a large-scale dataset.
@@ 7.2 Results
We observed comparable results for both tasks, and report results for Task1 in the main text, while the detailed results of Task2 are provided in Appendix B.
Quality vs. Compute Complexity    In Fig. 1, we depict the relationship between quality and compute complexity. The results show that Wukong consistently outperforms all baselines across various complexity levels, achieving over 0.2% improvement in LogLoss. Notably, Wukong holds its scaling law across two orders of magnitude in model complexity – approximately translating to a 0.1% improvement for every quadrupling of complexity. Among baselines, AFN+, DLRM and FinalMLP tend to reach a plateau after a certain complexity level, while AutoInt+, DCNv2 and MaskNet failed to further enhance quality 111 AutoInt+ and DCNv2 consistently faced significant training instability issue when further scaled up. AutoInt+ recovered from loss explosion, albeit with reduced model quality; while DCNv2 failed to recover, and its quality was estimated from performance before the explosion. MaskNet was hindered by excessive memory consumption, leading to out-of-memory errors, blocking further scaling up.. Nonetheless, even DCNv2, the top-performing baseline, demands a 40-fold increase in complexity to match Wukong’s quality.
Quality vs. Model Size    In Fig. 3, we illustrate the correlation between model quality and model size. Echoing the trends observed in compute complexity scaling above, Wukong consistently outperforms all baselines by rougly 0.2% across all scales of model size. while demonstrating a steady improvement trend up to over 637 billion parameters.
Figure 3: Scalability of Wukong with respect to # parameters.
Model-Specifc Scaling    Throughout the scaling process, we employed distinct strategies per model. Detailed hyperparameter settings for each run are provided in Appendix B. Scaling processes of each model are summarized as follows:
@@ 8 Discussions
Implication on Resource-Limited Research and Businesses    Wukong’s efficiency and scalability offer significant advantages to researchers and businesses with limited resources. Its capability to match the accuracy of other models with forty times less computing resources enables cutting-edge recommendation research and applications, even in environments with limited computation resources.
@@ Appendix A Model-Specific Grid Search Space on Criteo
Appendix B Model-Specific Scaling-up Configurations

### v4

@@ Abstract
Scaling laws play an instrumental role in the sustainable improvement in model quality. Unfortunately, recommendation models to date do not exhibit such laws similar to those observed in the domain of large language models, due to the inefficiencies of their upscaling mechanisms. This limitation poses significant challenges in adapting these models to increasingly more complex real-world datasets. In this paper, we propose an effective network architecture based purely on stacked factorization machines, and a synergistic upscaling strategy, collectively dubbed Wukong, to establish a scaling law in the domain of recommendation. Wukong’s unique design makes it possible to capture diverse, any-order of interactions simply through taller and wider layers. We conducted extensive evaluations on six public datasets, and our results demonstrate that Wukong consistently outperforms state-of-the-art models quality-wise. Further, we assessed Wukong’s scalability on an internal, large-scale dataset. The results show that Wukong retains its superiority in quality over state-of-the-art models, while holding the scaling law across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example, where prior arts fall short.
@@
Figure 1: Wukong outperforms existing state-of-the-art models while demonstrating a scaling law in the recommendation domain across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example.
@@ 1 Introduction
Deep learning-based recommendation systems (DLRS) power a wide range of online services today (Naumov et al., 2019; Wang et al., 2021a; Lian et al., 2021; Liu et al., 2022; Covington et al., 2016).
@@
To date, the primary trend of DLRS up-scaling is through sparse scaling, i.e., expanding the sizes of embedding tables (more rows and/or higher dimensions) for less collision and better expressiveness. Consequently, DLRS have reached trillions of parameters (Kang et al., 2020; Mudigere et al., 2021; Lian et al., 2021) with embedding tables dominating the parameter count. Unfortunately, this traditional way of up-scaling has a few practical drawbacks. Merely expanding the sparse component of a model does not enhance its ability to capture the complex interactions among an increasing number of features. Moreover, this trend notably diverges from the trend of hardware advancements, as most improvements in the next generation accelerators lie in the compute capacity (Luo et al., 2018, 2017), which embedding table lookups cannot utilize. Thus, simply expanding embedding table leads to prohibitive infrastructure costs with suboptimal accelerator utilization, especially in distributed settings (Luo et al., 2024).
@@
Our work aims to find an alternative scaling mechanism for recommendation models, that can establish a scaling law, similar to that established in the LLM domain. Namely, we would like to devise a unified architecture whose quality can be continuously improved in conjunction with dataset size, compute and parameter budgets, with a synergistic strategy.
@@
We focus on upscaling interaction components, dubbed dense scaling, to mitigate the quality and efficiency drawbacks from sparse scaling. However, existing models cannot benefit from this paradigm for various reasons. For example, DLRM lacks the ability to capture <higher-order> interactions; DCNv2 and AutoInt+ lack strategy for effective upscaling, leading to rapidly diminishing returns when scaling up; further, even with modern tricks like residual connection (He et al., 2016), layernorm (Ba et al., 2016), gradient clip (Pascanu et al., 2013)), up-scaling existing models is prone to training stability issues (Tang et al., 2023).
@@
To establish a scaling law for recommendation models, we propose Wukong, a simple interaction architecture that exhibits effective dense scaling properties. Inspired by the principles of binary exponentiation, our key innovation is to use a series of stacked Factorization Machines (FMs) to efficiently and scalably capture any-order feature interactions. In our design, each FM is responsible of capturing second order interactions with respect to its inputs, and the outputs from these FMs are subsequently transformed by MLPs into new embeddings, which encode the interactions results and serve as inputs to the next layers.
@@
We evaluated Wukong’s performance using six public datasets and a large-scale internal dataset. The results demonstrate that Wukong outperforms state-of-the-art models across all public datasets in terms of AUC, indicating the effectiveness of Wukong’s architecture and its ability to generalize across a wide range of recommendation tasks and datasets. In our internal dataset, Wukong not only significantly outperforms existing models in terms of quality at comparable levels of complexity but also shows continuous enhancements in quality when scaled up across two orders of magnitude in model complexity, extending beyond 100 GFLOP/example, where prior arts fall short.
@@ 2 Related Work
Dense Interaction Architectures    Capturing interaction between features is the key to DLRS effectiveness, and we highlight some of the prior arts. AFN+ (Cheng et al., 2020) transforms features into a logarithmic space to capture arbitrary order of interactions; AutoInt+ (Song et al., 2019) uses multi-head self-attention; DLRM and DeepFM (Naumov et al., 2019; Guo et al., 2017) leverage Factorization Machines (FM) (Rendle, 2010) to explicitly capture second order interactions; HOFM (Blondel et al., 2016) optimizes FM to efficiently capture higher order of interactions; DCNv2 (Wang et al., 2021a) uses CrossNet, which captures interactions via stacked feature crossing, which can be viewed as a form of elementwise input attention. FinalMLP (Mao et al., 2023) employs a bilinear fusion to aggregate results from two MLP streams, each takes stream-specific gated features as input. MaskNet (Wang et al., 2021b) adopts a series of MaskBlocks for interaction capture, applying “input attention” to the input itself and intermediate activations of DNN; xDeepFM (Lian et al., 2018) combines a DNN with a Compressed Interaction Network, which captures interactions through outer products and compressing the results with element-wise summation.
Scaling up DLRS     (Kang et al., 2020; Mudigere et al., 2021; Lian et al., 2021) provides mechanisms on sparse scaling.  (Shin et al., 2023) focuses on scaling up user representation models, with the largest model reported having a total compute less than 0.1 PF-days,  (Zhang et al., 2023) aims to improve sequence modeling on the user side, with the largest model reported having less than 0.8B parameters. Additionally  (Ardalani et al., 2022) studied the scaling law of DLRM, which is incorporated as a baseline in our work and further scaled up in our experiments. Orthogonally, (Zhao et al., 2023b) proposes a user-centric ranking formulation to improve scalability;  (Guo et al., 2023) provided insights on sparse scaling, demonstrating limits on prior arts and is complement to our work. Further, VIP5 (Geng et al., 2023) leverages existing scaling laws in LLMs to apply a multimodal LLM to recommendation, however,  (Lin et al., 2023) points out that further study is needed to verify whether larger implies better in LLM-powered recommenders, while  (Huang et al., 2024) suggests evaluations on more diverse datasets are needed to for a conclusion.
@@ 3.2 Embedding Layer
Given a multi-hot categorical input, an embedding table maps it to a dense embedding. This process involves a series of lookups, each corresponding to a “hot” dimensions within the input. The lookup results are then aggregated using a pooling operation (usually a summation).
In our design, the embedding dimension is standardized for all embeddings generated by the Embedding Layer, known as the global embedding dimension $d$. To accommodate the varying significance of different features, multiple embeddings are generated for each feature that is deemed significant. Less important features are allocated smaller underlying embedding dimensions. These smaller embeddings are then collectively grouped, concatenated, and transformed into $d$-dimensional embeddings using a MLP.
@@ 3.3 Interaction Stack
The interaction modules stack $l$ identical interaction layers, where each layer captures progressively higher-order feature interactions using Factorization Machines (FMs).
@@ 3.4 Factorization Machine Block (FMB)
A FMB contains a FM followed by a MLP. The FM is used to capture explicit feature interactions of the input embeddings, with the output being a 2D interaction matrix where each element represents the interaction between a pair of embeddings. This interaction matrix is flattened and converted to a vector with shape of $(n_{F}\times d)$ through the MLP, and reshaped to $n_{F}$ embeddings for later use.
@@ 3.6 Optimized FM
FM’s computation and storage complexity grows quadratically with the number of embeddings with the pair-wise dot product, and This quickly becomes prohibitive on real-world datasets with thousands of features.
@@ 3.8 Scaling Wukong
We now summarize the main hyperparameters that are related to scale up and later we describe our efforts to upscaling Wukong with respect to these hyperparameters.
@@ 3.9 Intuition Behind Wukong’s Enhanced Effectiveness
Additionally, Wukong treats each embedding as a single unit, focusing on embedding-wise interactions. This approach significantly reduces computational demands compared to architectures that capture element-wise interactions.
@@ 4 Implementation
Overall, distributed training is required to make Wukong training feasible. For the embedding layer, we use a column-wise sharded embedding bag implementation provided by Neo (Mudigere et al., 2021) and NeuroShard (Zha et al., 2023). On the dense part, we balance the trade-off between performance and memory capacity by adopting FSDP (Zhao et al., 2023a) and tune the sharding factor so that the model fits in the memory without creating too much redundancy.
To enhance training efficiency, we employ automatic operator fusion through to improve training performance. In addition, we aggressively apply quantization to reduce compute, memory, and communication overheads simultaneously. Specifically, we train Wukong’s embedding tables in FP16, and communicate embedding lookup results in FP16 in the forward pass and BF16 in the backward pass; we use BF16 quantization during the transport gradients for dense parameters in the backward pass.
@@ 6.1.3 Metrics
AUC    Area Under the Curve (AUC) measures the model’s ability to correctly classify positives and negatives across all thresholds. Higher the better. We use AUC as the basis for hyperparameter tuning and topline metric for reporting, following recommendation conventions (Tien et al., 2014; Blondel et al., 2016; Song et al., 2019; Wang et al., 2021a; Zhu et al., 2022b; Mao et al., 2023).
LogLoss    The log loss quantifies the penalty based on how far the prediction is from the actual label. Lower the better.
@@ 7 Evaluation on an Internal Dataset
In this section, we show the scalability of Wukong and gain a deep understanding of how different individual components of Wukong contribute to its effectiveness, using a large-scale dataset which enables the study for merging properties that is not seen in small, public datasets.
@@ 7.2 Results
We observed comparable results for both tasks, and report results for Task1 in the main text, while the detailed results of Task2 are provided in Appendix C.
Quality vs. Compute Complexity    In Fig. 1, we depict the relationship between quality and compute complexity (empirically, $y=-100+99.56x^{0.00071}$). The results show that Wukong consistently outperforms all baselines across various complexity levels, achieving over 0.2% improvement in LogLoss. Notably, Wukong holds its scaling law across two orders of magnitude in model complexity – approximately translating to a 0.1% improvement for every quadrupling of complexity. Among baselines, AFN+, DLRM and FinalMLP tend to reach a plateau after a certain complexity level, while AutoInt+, DCNv2 and MaskNet failed to further enhance quality 111 AutoInt+ and DCNv2 consistently faced significant training instability issue when further scaled up. AutoInt+ recovered from loss explosion, albeit with reduced model quality; while DCNv2 failed to recover, and its quality was estimated from performance before the explosion. MaskNet was hindered by excessive memory consumption, leading to out-of-memory errors, blocking further scaling up.. Nonetheless, even DCNv2, the top-performing baseline, demands a 40-fold increase in complexity to match Wukong’s quality.
Quality vs. Model Size    In Fig. 3, we illustrate the correlation between model quality and model size. Echoing the trends observed in compute complexity scaling above, Wukong consistently outperforms all baselines by roughly 0.2% across all scales of model size. while demonstrating a steady improvement trend up to over 637 billion parameters [^2] We verified Wukong’s effectiveness in both online and offline settings, and for brevity, we focus on reporting offline metrics.
Quality vs. Data Size    See Appendix E.
Figure 3: Scalability of Wukong with respect to # parameters on the internal dataset.
Model-Specifc Scaling    Throughout the scaling process, we employed distinct strategies per model. Detailed hyperparameter settings for each run are provided in Appendix C. Scaling processes of each model are summarized as follows:
@@ Appendix A Model-Specific Grid Search Space on Criteo
Appendix B Model Complexity/Size on Public Datasets
Please refer to Table 3 for details.
[…]
      AFN+ 13607961 81.45 3205771 10.57 11259899 33.36
AutoInt+ 306151 15.98 1989378 10.74 1293590 3.00
DCNv2 7251073 39.35 3454209 11.84 1243101 2.02
DLRM 1527073 4.99 2853761 12.66 1238421 1.99
FinalMLP 3115954 10.66 3020498 9.25 1040902 0.81
MaskNet 21189593 122.92 2053249 7.87 2624571 10.30
xDeepFM 66206 0.01 5147446 5.91 994439 0.00
Wukong 7769589 44.09 2219409 9.08 17734369 37.49
[…]
      AFN+ 84013143 10.95 167888773 13.45 26436527675 1826.50
AutoInt+ 43937794 79.18 167395330 28.41 26142619137 163.34
DCNv2 42636609 9.10 193472513 166.19 26179364097 262.35
DLRM 41446513 1.99 42192033 4.53 26136307201 4.62
FinalMLP 580227666 12.17 85624114 16.39 26149500924 84.95
MaskNet 41833034 4.29 42353201 5.44 26160209153 147.40
xDeepFM 53912381 6.81 168867028 2.98 26300221171 41.57
Wukong 44671649 22.27 175724173 63.00 26163636001 173.73
Table 3: Model complexity and size on public datasets.
Appendix C Model-Specific Scaling-up Configurations
Please refer to Table 5 for details.
Appendix D Analysis of High Order Interactions in Wukong
$\min\underset{i,j\in S}{\Sigma}r_{ij}-{X^{1}}{X^{1}}^{T}$
Appendix E Scaling Law on Training Data Volume
Figure 6: Wukong’s model quality improvements versus training data volume and training compute.
Fig. 6 provides a summary for Wukong’s performance versus the dataset size on which it is trained on (one pass). Similar to what has been observed on LLMs, we found that large models are more data-efficient, meaning that they require fewer samples to achieve the same quality improvement. In addition, we found that all Wukong models have consistently improved their model quality up to the end of 146B data, while larger models have a steeper trend in the model quality improvement. We also noticed that one of the limitations of our study is the dataset size is still far less for the large model to converge, which will be one of the areas for further study.
Appendix F Comparing with Transformer-based Approaches
We highlight differences and provide intuitions on why Wukong scales better than Transformer-based approaches like AutoInt+ (Song et al., 2019). While the structure of Wukong resembles that of Transformer’s, we note the following architectural difference: first, the projection used in Wukong is MLP (bit-wise) in both FMB and each layer instead of FFN (embedding/position-wise) in transformer; second, Wukong is configured as a pyramid shape versus the uniform shape used in transformers.
We hypothesize that the difference in projection plays an important role in quality delivery. These MLPs operate over the flattened input embeddings, essentially providing each feature with a different projection matrix. Our intuition is that this helps the model learn from heterogeneous input features, contrasting to a single embedding space used in LLMs. Similar intuition was discussed in (Gui et al., 2023).
Efficiency-wise, we argue that the pyramid shape configuration allows Wukong to exclude unnecessary computations by contracting the number of embeddings used in each layer.
To verify these hypothesis, we conduct the following experiments by applying Wukong’s unique components to Autoint+, which conclude (1) using bit-wise MLP instead of FFN for V-projection improves LogLoss by 0.34%; (2) adding bit-wise MLPs after self-attentions improves LogLoss by 0.65%; (3) combining both, along with a pyramid layer shape (by using LCB on the first layer’s output) achieves 0.57% quality improvement. Compared to the scaled up Autoint+, Wukong achieves 0.08% quality improvement, while saves 90% FLOPs. We summarize the results in Table 4
   Changes Relative GFLOP
vs. Autoint+ LogLoss (%) /example
  Vanilla Autoint+ 0 8
V=FFN() $\to$ V=MLP() -0.34 13
Scaled-up Autoint+ -0.49 50
V=FFN() $\to$ V=MLP() -0.57 5
Layer FFN $\to$ Layer MLP
Pyramid layer shape
Layer FFN $\to$ Layer MLP -0.65 36
Table 4: Replacing/Adding Wukong’s unique components to vanilla Autoint+ improves the model quality.